# Machine learning classification of black holes in the mass–spin diagram

### v2

@@ III Classifying black holes
Machine learning algorithms require comprehensive datasets for training otherwise the classification predictions degrade. This can happen in at least two main ways which limit their ability to generalize predictions: either the dataset sparsely fills the total parameter space volume of the system, or the dataset exhibits degeneracies such as overlapping contributions from multiple components. In this section, we will explore both cases (i.e., the former in Subsec. III.1, and the latter in Subsec. III.2) in the context of black-hole binary populations.
@@ III.1 The Black Hole Continuum
Intermediate-mass black holes (IMBHs; $\approx 10^{2}\,\rm\,M_{\odot}$ to $10^{5}\,\rm\,M_{\odot}$) occupy a contentious region between well-established stellar-mass and supermassive regimes. EM candidates in ultra-luminous and hyper-luminous X-ray sources, dynamical globular clusters, or dwarf galaxies are numerous but have yet to be confirmed confidently [21]. Meanwhile, recent GW events with LVK (e.g., GW190521, GW200220, GW231123) concretely confirm the presence of lower-end IMBHs, offering glimpses into this apparent mass gap.
The traditional tripartite division—stellar, intermediate, supermassive—is a human-constructed taxonomy based on discovery methods and observational biases: EM surveys select high-luminosity or high-mass remnants, and GW detections sample different redshift and mass ranges. For example, pulsar timing arrays are more sensitive to high-mass, low redshift supermassive black-hole binaries, and LISA will be sensitive to lower mass black holes to high redshift. These selection effects correlate with mass–spin trends inherently tied to astrophysical origins. In this subsection we examine whether machine learning with Scikit-learn [41] can discover these human-assigned distinctions using only mass and spin features.
First, we generate a toy dataset comprising two populations: stellar-mass black holes with spin magnitudes uniformly distributed between 0 and 1, and supermassive black holes with spins uniformly between 0.9 and 1. Figure 5 shows how a Gaussian mixture model (left panel) and k-means (right panel) for clustering on this bimodal distribution successfully separates stellar and supermassive populations due to the simple mass-spin correlations. We perform this clustering in log-mass space to ensure proper feature scaling. This hints toward a potential for machine learning methods to reproduce—and ultimately refine—existing classifications, anticipate new categories as datasets grow, and identify selection biases inherent in human-based taxonomy.
@@
Spectral clustering (with nearest neighbor affinity), which operates on the full mass–spin manifold, can struggle to accurately recover the tripartite division, as shown by the three color-marker pairs in Fig. 6 where the vertical black lines indicate the true mass boundaries assumed in constructing the theoretical dataset. The misclassification of edge cases reflects the spin distinctions we imposed, but more importantly it demonstrates that such algorithms can detect latent structure in complex data which may arise from selection effects or data analysis assumptions. Also, it provides further motivation for the development of a black hole continuum mass function, which is an active area of research [111] and would present an interesting challenge for these strategies. We note that this classification was performed without latent space embedding prior to clustering, implying that an analysis which includes latent space transformation may perform even better, which we leave to future work.
@@ III.2 Deep Learning of Stellar-mass Black Holes
In case 1., we take the data plotted in Fig. 4 by taking the secondary mass $m_{2}$ and spin magnitudes $\chi_{2}$ from the Bavera et al. 2020 and Steinle & Kesden (2021) models as these produce the least overlapping regions in Fig. 4. We limit spin magnitudes to between $10^{-4}$ and $1.0$ to remove non-rotating systems, apply a log transform, and standardize the masses and spins separately via a Gaussian offset $\mathcal{N}(\log_{10}\leavevmode\nobreak\ x,0.01)$ where $x$ is the mass or spin, and pairing each black-hole binary with corresponding true model labels. The standardized mass and spin features are combined into two-dimensional feature vectors $(\mathrm{m},\mathrm{\chi})$ for each sample with corresponding integer true class labels for each model, producing the final feature–label (i.e., input–output) datasets.
@@
Using PyTorch [40] we implement and benchmark a neural network architecture which imposes a regularized latent structure, and we perform clustering with scikit-learn [41]. We do this in a modular, end-to-end machine learning framework to extract latent structure in simulated black hole populations and perform classification  [^4] Codes used in this study will be made publicly available upon publication. The workflow follows a standard pipeline, where raw simulation outputs are extracted, processed, embedded into a compressed latent space with PyTorch neural net architecture, and clustered with the scikit-learn machine learning suite. We use an unsupervised classification scheme via a variational autoencoder (VAE) and k-means clustering, extending it to a semi-supervised classification framework, and a random forest supervised classifier. Raw data are split into 60% training, 20% test, and 20% validation sets containing black hole masses, spin magnitudes, and true class labels. As the performance is the same to within 1% between the training, test, and validation datasets, we only show the training results for brevity. All three datasets contain 100,000 black holes in each class, but results in Table 1 are similar for 10,000 black holes each.
@@
The classification and benchmarking results for the three datasets are shown in Table 1, where we only report the training results for brevity as the test and validation results are similar (i.e. $\lesssim 10$% difference). The unsupervised k-means clustering is unable to correctly learn the highly overlapping mass–spin distributions. This is demonstrated by the $\approx 0$ ARI and NMI indicating ineffective latent space learning. The consistent F1-score of $\sim 0.5$ for k-means across different hyperparameter settings combined with the near-0 ARI and NMI indicates that the k-means approach ends up with a random $\sim 50/50$ clustering assignment. We also confirm independently that the ARI and NMI are 0 ($\sim 0.3$) when unsupervised (supervised) clustering is performed directly on the raw input data rather than on the latent representation space. We also performed this analysis with a Gaussian mixture model instead of k-means and it performed similarly.
Figure 7 demonstrates the difference in clustering in the trained VAE latent space between the unsupervised (top row panels) and supervised (bottom row panels) clustering of dataset 1., where the true and predicted data correspond to the left and right columns, respectively. Blue circles and orange squares correspond to the two spin models of Bavera et al. (2020) and Steinle & Kesden (2021), respectively. This visually confirms the challenge of latent space learning for the unsupervised approach. The success of the supervised random forest classifier (i.e.,  98% accuracy) is shown in the bottom-row panels of Fig. 7, as the two predicted clusters are closely reproduced in the right panel with respect to the true clusters in the left panel. The dataset 1. has more distinct classes compared to datasets 2. and 3., and the spins of the two classes in dataset 1. correspond to the same black hole masses but in datasets 2. and 3. the spins of the two classes have different corresponding masses; both of these features contribute to the random forest performing better on dataset 1.
@@
While the unsupervised approach will require further investigation in future work, the success of the supervised random forest classifier indicates that a semi-supervised approach might show intermediate performance. Thus, we expanded the k-means pipeline for semi-supervised classification in which we assume some fraction has known labels and the training for the unlabeled fraction is the same as in the unsupervised pipeline but training for the labeled fraction employs an additional penalty in the loss function where points farther away from each other in the parameter space have less weight via cross-entropy loss in PyTorch. The result of the semi-supervised approach is sensitive to the initialization of random states in several components of the pipeline. , so we ran the semi-supervised k-means pipeline with 5 different initializations and a labeled fraction of 0.5. The resultant best semi-supervised clustering had accuracy $\approx$ 0.6, f1-score $\approx$ 0.58, and NMI and ARI $\approx$ 0.45. Comparing with the fully unsupervised case, this demonstrates that the low-dimensional latent space transformation can improve clusterability, but a full analysis is beyond the scope of the present study.
The use of uncertainty in the event parameters is a key area of future development, suggesting the use of methods like Bayesian Neural Networks that can quantify uncertainty in training and prediction. Such an analysis can be paired with hierarchical inference output [29], allowing the ML algorithms to operate directly on the posteriors rather than summary statistics (i.e. medians), enabling better utilization of astrophysical information for latent space learning. Our results indicate that combining the mass–spin diagram with machine learning contains potential for many applications in studying astrophysical black hole populations where distinguishing formation channels which is a great open problem in gravitational-wave astronomy, see e.g.’s [114, 115, 116, 117, 55, 118, 119]. This motivates development of new unsupervised and semi-supervised learning methods, and fine-tuning of the approaches attempted here, for challenging astrophysical datasets from future GW and EM detectors, which we leave to a followup study [120]. Another avenue of future work will use more realistic combined populations of black-hole binary formation channels in a cosmological setting, i.e., the redshift evolution across the mass–spin diagram.
@@ IV Summary
First we examined the mass–spin diagram in the context of the three black hole mass regimes: stellar-mass black holes formed from gravitational stellar core collapse, intermediate-mass black holes possibly formed through runaway collisions or direct collapse, and supermassive black holes from light and heavy seeds. We approximate redshift evolution using physically motivated mass and spin prescriptions, incorporating both super-Eddington accretion and stochastic processes. These black hole trajectories serve as tracks in the mass–spin space, shaped by their accretion history and seed properties. For instance, stellar-mass BHs can spin up efficiently via fallback or disk accretion, while intermediate-mass black holes undergo diverse accretion and merger episodes, and supermassive black holes experience coherent spin-up in the early Universe and hierarchical galaxy mergers. The mass–spin trajectories carry a fossil record of evolutionary mechanisms, analogous to stellar tracks and clusters in the HR diagram, motivating inference of a black hole ’main sequence’ driven by accretion through cosmic evolution.
@@ V Discussion
Out results have multiple implications. First the mass–spin diagram provides a visual and statistical framework for classifying BHs across all mass scales, allowing direct comparison of theory and data in a physically meaningful space. Machine learning of gravitational-wave populations in the diagram is useful for studying formation possibilities. This framework can be improved by incorporating merger products, eccentricity, environmental gas effects, and host galaxy priors into the mass–spin space. As observed catalogs grow in size and fidelity, these methods will become increasingly useful for inferring formation histories of black holes. In the HR diagram, stars on the main sequence occupy the primary band defined by core hydrogen burning in hydrostatic equilibrium, a fundamental and long-lived phase. We suggest that analogues for black holes may be tied to mass–spin growth tracks, such as prolonged, coherent accretion episodes, as the black hole evolves in redshift. Numerous theoretical and cosmological simulation studies support the universal nature of accretion-driven mass and spin growth [123, 124, 125, 126]. These simulations imply that gas accretion dominates supermassive black hole growth in the high-redshift universe allowing for efficient spin-up, while galaxy mergers dominate at lower redshifts, possibly providing random spin orientations relative to the host galaxy; together, these create mass–spin tracks for intermediate and supermassive black holes. A recent study found a three-phase evolution, depending on the mass and redshift of the black hole, with predictions for their spin magnitudes and orientations relative to their host galaxies [127]. General relativistic magneto-hydrodynamic simulations show the interplay of accretion physics and generic spin evolution in gravitational-wave events [128].
@@
A “Cosmic Accretion” main sequence serves as a locus in mass–spin space determined by accretion physics across diverse formation environments and pathways. Deviations from a track, such as low-spin natal black holes, hierarchical merger remnants, or chaotic accretors, can play the role of post-main-sequence or blue-loop outliers. Black hole binaries in isolation will be static on the diagram unless they merge, resulting in a merger-tree structure analogue to the main sequence, where black holes appear and disappear as merger products or remnants. Ultimately, the many main sequences would be connected through cosmological evolution of the Universe.
@@
The future of GW astronomy is bright with the construction and planning of many detectors across the GW frequency spectrum, with the Square Kilometer array [129], Meerkat [130], and DAS-2000 [131] in the nano-Hz regime, astrometric facilities such as Theia [132] in the micro-Hz regime, LISA [65], TianQin [133], and Taiji [134] in the milli-Hz regime, and Cosmic Explorers [121] and the Einstein Telescope [122] in the Hz to kilo-Hz regime. Interestingly, one can perform similar analysis as done here on a diagram defined by GW-frequency and black hole spin to compare the spins of binary mergers as a function of GW frequency. Multi-band studies can compare the populations detectable by pulsar timing arrays and LISA (supermassive black holes) with those detectable by LVK, Cosmic Explorer and Einstein Telescope (stellar-mass black holes), by their spin evolution and with both sets of detectors peering into the intermediate-mass black hole regime. Ultimately, the potential for multi-messenger applications of the mass–spin diagram are also plentiful with the development of future X-ray facilities such as AXIS [135] which will provide unprecedented measurements of accreting systems. The classification of joint EM and GW populations, as done in Fig. 6, will be an important direction of future work.

### v4

@@ III Classifying black holes
Machine learning algorithms require comprehensive datasets for training otherwise the classification predictions degrade. This can happen in at least two main ways which limit their ability to generalize predictions: either the dataset sparsely fills the total parameter space volume of the system, or the dataset exhibits degeneracies such as overlapping contributions from multiple components. In this section, we will explore both cases (i.e., the former in Subsec. III.1, and the latter in Subsec. III.2) in the context of black-hole binary populations. We construct these datasets from toy-model data or from the raw output of realistic simulations in order to test this machine learning framework in a controlled setting, and we leave application to real-world datasets (i.e. from gravitational and electromagnetic observations) to a future study.
@@ III.1 The Black Hole Continuum
Intermediate‑mass black holes (IMBHs; $\approx 10^{2}\,\rm\,M_{\odot}$ to $10^{5}\,\rm\,M_{\odot}$) occupy a contentious region between well-established stellar‑mass and supermassive regimes. EM candidates in ultra‑luminous and hyper-luminous X‑ray sources, dynamical globular clusters, or dwarf galaxies are numerous but have yet to be confirmed confidently [23]. Meanwhile, recent GW events with LVK (e.g., GW190521, GW200220, GW231123) concretely confirm the presence of lower-end IMBHs, offering glimpses into this apparent mass gap.
The traditional tripartite division—stellar, intermediate, supermassive—is a human‑constructed taxonomy based on discovery methods and observational biases: EM surveys select high‑luminosity or high‑mass remnants, and GW detections of stellar-mass and intermediate mass black holes sample different redshift and mass ranges. For example, pulsar timing arrays are more sensitive to high-mass, low redshift supermassive black-hole binaries, and LISA will be sensitive to lower mass black holes to high redshift. These selection effects correlate with mass–spin trends inherently tied to astrophysical origins. In this subsection we examine whether machine learning with Scikit-learn [45] can discover these human‑assigned distinctions using only mass and spin features.
First, we generate a toy dataset comprising two populations that reflect observed masses and spins (see Table 1): stellar‑mass black holes with spin magnitudes uniformly distributed between 0 and 1, and supermassive black holes with spins uniformly between 0.9 and 1. Figure 5 shows how a Gaussian mixture model (left panel) and k-means (right panel) for clustering on this bimodal distribution successfully separates stellar and supermassive populations due to the simple mass-spin correlations. We perform this clustering in log-mass space to ensure proper feature scaling. This hints toward a potential for machine learning methods to reproduce—and ultimately refine—existing classifications, anticipate new categories as datasets grow, and identify biases inherent in human-based taxonomy.
@@
Spectral clustering (with nearest neighbor affinity), which operates on the full mass–spin manifold, can struggle to accurately recover the tripartite division, as shown by the three color-marker pairs in Fig. 6 where the vertical black lines indicate the true mass boundaries assumed in constructing the theoretical dataset. The misclassification of edge cases reflects the spin distinctions we imposed, but more importantly it demonstrates that such algorithms can detect latent structure in complex data which may arise from selection effects or data analysis assumptions. Also, it provides further motivation for the development of a black hole continuum mass function, which is an active area of research [114] and would present an interesting challenge for these strategies. While the IMBHs are composed of two subclasses with respect to the spin axis, the algorithm did not identify them as separate clusters, likely due to our assumption of three classes in the learning. In future work, we will explore how this depends on the number of classes one assumes in the unsupervised learning and its implications for future observations of IMBHs. We note that this classification was performed without latent space embedding prior to clustering, implying that an analysis which includes latent space transformation may perform even better, which we leave to future work.
@@ III.2 Deep Learning of Stellar-mass Black Holes
In case 1., we take the data plotted in Fig. 4 by taking the secondary mass $m_{2}$ and spin magnitudes $\chi_{2}$ from the Bavera et al. 2020 and Steinle & Kesden (2021) models as these produce the least overlapping regions in Fig. 4. We limit spin magnitudes to between $10^{-4}$ and $1.0$ to remove non-rotating systems, apply a log transform, and standardize the masses and spins separately via a Gaussian offset $\mathcal{N}(\log_{10}\penalty 10000\ x,0.01)$ where $x$ is the mass or spin, and pairing each black-hole binary with corresponding true model labels. The standardized mass and spin features are combined into two-dimensional feature vectors $(\mathrm{m},\mathrm{\chi})$ for each sample with corresponding integer true class labels for each model, producing the final feature–label (i.e., input–output) datasets.
@@
Using PyTorch [44] we implement and benchmark a neural network architecture which imposes a regularized latent structure, and we perform clustering with scikit-learn [45]. We do this in a modular, end-to-end machine learning framework to extract latent structure in simulated black hole populations and perform classification  [^5] Codes used in this study will be made publicly available upon publication. The workflow follows a standard pipeline, where raw simulation outputs are extracted, processed, embedded into a compressed latent space with PyTorch neural net architecture, and clustered with the scikit-learn machine learning suite. We use an unsupervised classification scheme via a variational autoencoder (VAE) and k-means clustering, extending it to a semi-supervised classification framework, and a random forest supervised classifier. Raw data are split into 60% training, 20% test, and 20% validation sets containing black hole masses, spin magnitudes, and true class labels. As the performance is the same to within 10% between the training, test, and validation datasets, we only show the training results for brevity. All three datasets contain 100,000 black holes in each class, but results in Table 2 are similar for 10,000 black holes each.
@@
The classification and benchmarking results for the three datasets are shown in Table 2, where we only report the training results for brevity as the test and validation results are similar (i.e. $\lesssim 10$% difference). The unsupervised k-means clustering is unable to correctly learn the highly overlapping mass–spin distributions. This is demonstrated by the $\approx 0$ ARI and NMI indicating ineffective latent space learning. The consistent F1-score of $\sim 0.5$ for k-means across different hyperparameter settings combined with the near-0 ARI and NMI indicates that the k-means approach ends up with a random $\sim 50/50$ clustering assignment. We also confirm independently that the ARI and NMI are 0 ($\sim 0.3$) when unsupervised (supervised) clustering is performed directly on the transformed raw input data rather than on the latent representation space. We also performed this analysis with a Gaussian mixture model instead of k-means and it performed similarly.
Figure 7 demonstrates the difference in clustering in the trained VAE latent space between the unsupervised (top row panels) and supervised (bottom row panels) clustering of dataset 1., where the true and predicted data correspond to the left and right columns, respectively. Blue circles and orange squares correspond to the two spin models of Bavera et al. (2020) and Steinle & Kesden (2021), respectively. This visually confirms the challenge of latent space learning for the unsupervised approach. The success of the supervised random forest classifier (i.e., $\sim 90\%$ accuracy) is shown in the bottom-row panels of Fig. 7, as the two predicted clusters are closely reproduced in the right panel with respect to the true clusters in the left panel. The dataset 1. has more distinct classes compared to datasets 2. and 3., and the spins of the two classes in dataset 1. correspond to the same black hole masses but in datasets 2. and 3. the spins of the two classes have different corresponding masses; both of these features contribute to the random forest performing better on dataset 1.
@@
While the unsupervised approach will require further investigation in future work, the success of the supervised random forest classifier indicates that a semi-supervised approach might show intermediate performance. Thus, we expanded the k-means pipeline for semi-supervised classification in which we assume some fraction has known labels and the training for the unlabeled fraction is the same as in the unsupervised pipeline but training for the labeled fraction employs an additional penalty in the loss function where points farther away from each other in the parameter space have less weight via cross-entropy loss in PyTorch. The result of the semi-supervised approach is sensitive to the initialization of random states in several components of the pipeline. We ran the semi-supervised k-means pipeline with 5 different initializations and a labeled fraction of 0.5. The resultant best semi-supervised clustering had accuracy $\approx$ 0.6, f1-score $\approx$ 0.58, and NMI and ARI $\approx$ 0.45. Comparing with the fully unsupervised case, this demonstrates that the low-dimensional latent space transformation can improve clusterability, but a full analysis is beyond the scope of the present study.
The use of uncertainty in the event parameters is a key area of future development, suggesting the use of methods like Bayesian Neural Networks that can quantify uncertainty in training and prediction. Such an analysis can be paired with hierarchical inference output [33, 117], allowing the ML algorithms to operate directly on the posteriors rather than summary statistics (i.e. medians), enabling better utilization of astrophysical information for latent space learning. Our results indicate that combining the mass–spin diagram with machine learning contains potential for many applications in studying astrophysical black hole populations where distinguishing formation channels which is a great open problem in gravitational-wave astronomy, see as examples [118, 119, 120, 121, 59, 122, 123]. This motivates development of new unsupervised and semi-supervised learning methods, and fine-tuning of the approaches attempted here, for challenging astrophysical datasets from future GW and EM detectors, which we leave to a followup study [124]. Another avenue of future work will use more realistic combined populations of black-hole binary formation channels in a cosmological setting, i.e., the redshift evolution across the mass–spin diagram.
@@ IV Summary
First we examined the mass–spin diagram in the context of the three black hole mass regimes: stellar-mass black holes formed from gravitational stellar core collapse, intermediate-mass black holes possibly formed through runaway collisions or direct collapse, and supermassive black holes from light and heavy seeds. We approximate redshift evolution using physically motivated mass and spin prescriptions, incorporating both super-Eddington accretion and stochastic processes. These black hole trajectories serve as tracks in the mass–spin space, shaped by their accretion history and seed properties. For instance, stellar-mass BHs can spin up efficiently via fallback or disk accretion, while intermediate-mass black holes undergo diverse accretion and merger episodes, and supermassive black holes experience coherent spin-up in the early Universe and hierarchical galaxy mergers. The mass–spin trajectories carry a fossil record of evolutionary mechanisms, analogous to stellar tracks and clusters in the HR diagram.
@@
In this work, the number of clusters in the unsupervised classification of Fig. 6 was fixed to three to reflect the canonical stellar-mass, intermediate-mass, and supermassive regimes and to test whether the algorithm could recover these established divisions. More generally, however, the optimal number of clusters is not known a priori in unsupervised learning and can be estimated through data-driven criteria such as the elbow method, silhouette score, or inspection of the Laplacian eigenspectrum gap in spectral clustering. Future applications of this framework to observational datasets will employ such methods to algorithmically infer the population structure. Likewise, the dimensionality of the variational autoencoder latent space, as shown in Fig. 7 and which was set to two for direct comparison with the mass–spin diagram, is related to the dimensionality of the mass–spin diagram and can be treated as a tunable hyperparameter. Exploring higher-dimensional latent representations, including additional astrophysical parameters such as spin orientation, redshift, or accretion rate, will form part of future work aimed at improving cluster separability and physical interpretability.
@@ V Discussion
Out results have multiple implications. First the mass–spin diagram provides a visual and statistical framework for classifying BHs across all mass scales, allowing direct comparison of theory and data in a physically meaningful space. Machine learning of gravitational-wave populations in the diagram is useful for studying formation possibilities. This framework can be improved by incorporating merger products, eccentricity, environmental gas effects, and host galaxy priors into the mass–spin space. As observed catalogs grow in size and fidelity, these methods will become increasingly useful for inferring formation histories of black holes. In the HR diagram, stars on the main sequence occupy the primary band defined by core hydrogen burning in hydrostatic equilibrium, a fundamental and long‐lived phase. We suggest that analogues for black holes may be tied to mass–spin growth tracks, such as prolonged, coherent accretion episodes, as the black hole evolves in redshift. Numerous theoretical and cosmological simulation studies support the universal nature of accretion-driven mass and spin growth [127, 128, 129, 130]. These simulations imply that gas accretion dominates supermassive black hole growth in the high-redshift universe allowing for efficient spin-up, while galaxy mergers dominate at lower redshifts, possibly providing random spin orientations relative to the host galaxy; together, these create mass–spin tracks for intermediate and supermassive black holes. A recent study found a three-phase evolution, depending on the mass and redshift of the black hole, with predictions for their spin magnitudes and orientations relative to their host galaxies [131]. General relativistic magneto-hydrodynamic simulations show the interplay of accretion physics and generic spin evolution in gravitational-wave events [132].
@@
Analogues to the stellar main sequence, which is the band in the HR diagram where stars spend the most time, can be found in the mass–spin diagram. An example we explore is Cosmic Accretion, a main sequence due to prolonged episodes of accretion across diverse formation environments and pathways, seen as mass–spin snapshots of all black holes within a certain redshift bin. Determining these snapshots accurately would require realistic models of the origin and evolution of black hole populations across mass and redshift —for stellar-mass black holes this is tied to metallicity-dependent star formation history and for more massive black holes is tied to the cosmological evolution of galaxy clusters. Mechanisms that drive mass and spin evolution possibilities of black holes create tracks through the mass–spin diagram which can be marked by continuous or discontinuous transitions, such as from accretion or mergers, respectively. Deviations from these tracks, such as mechanisms like low-spin natal black holes, hierarchical mergers, or chaotic accretors, can play the role of post-main-sequence stars or blue-loop outliers. Black hole binaries in isolation will be static on the diagram unless they merge, resulting in a merger-tree structure analogue type of main sequence, where black holes appear and disappear as merger products or remnants and tend to pile-up in regions of the diagram according to merger-dependent mass and spin distributions. Ultimately, the many main sequences would be connected through cosmological evolution of the Universe.
Our study yields three main new testable predictions:
  • the existence of mass–spin gaps in the diagram resulting from mass–spin correlations.
  • the existence of black hole main sequences tied to formation channels; channels with accretion will have distinct motions on the mass–spin diagram compared to channels with mergers.
  • the application of machine learning to classify black hole populations within and across mass regimes according to observed and theoretical spin prescriptions and correlations.
For example, in the stellar-mass regime, it is known that EM and GW observed populations differ in metallicity and redshift [111, 112], suggesting that machine learning classification can aid in categorizing and generalizing information of these distinctions from future datasets. There are important caveats to our analysis. We use various theoretical datasets to explore this new application of machine learning, but the application to real data will encounter new challenges not revealed in the present study. These datasets range from toy models of theoretical expectations and of observed correlations in real data to raw simulation output of realistic models. These products of theoretical formation channels of black holes play a crucial role in constructing useful training datasets for the machine learning algorithms. Also, we did not explore the motion of black holes across the diagram in our machine learning analysis - utilizing such black hole tracks will open a new area of application of the techniques demonstrated here.
@@
The future of GW astronomy is bright with the construction and planning of many detectors across the GW frequency spectrum, with the Square Kilometer array [133], MeerKAT [134], and DAS-2000 [135] in the nano-Hz regime, astrometric facilities such as Theia [136] in the micro-Hz regime, LISA [68], TianQin [137], and Taiji [138] in the milli-Hz regime, and Cosmic Explorers [125] and the Einstein Telescope [126] in the Hz to kilo-Hz regime. Interestingly, one can perform similar analysis as done here on a diagram defined by GW-frequency and black hole spin to compare the spins of binary mergers as a function of GW frequency. Multi-band studies can compare the populations detectable by pulsar timing arrays and LISA (supermassive black holes) with those detectable by LVK, Cosmic Explorer and Einstein Telescope (stellar-mass black holes), by their spin evolution and with both sets of detectors peering into the intermediate-mass black hole regime. Ultimately, the potential for multi-messenger applications of the mass–spin diagram are also plentiful with the development of future X-ray facilities such as AXIS [139] which will provide unprecedented measurements of accreting systems. The classification of joint EM and GW populations, as done in Fig. 6, will be an important direction of future work in multi-messenger astronomy.